# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ E.1 Background on optimal experimental design
The Kiefer-Wolfowitz theorem, see [Kiefer and Wolfowitz, 1960], shows that a D-optimal design also ensures that $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\left({\sum_{x\in\mathcal{X}}\pi(x)xx^{\top}}\right)^{-1}}^{2}=d$. A direct consequence is that $\mathcal{X}$ is a subset of the ellipsoid $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$. The Kiefer-Wolfowitz theorem can be interpreted by saying that the set $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$ is the minimum volume ellipsoid containing $\mathcal{X}$, see [Lattimore and Szepesvári, 2020, Theorem 21.1].
In the case where we have a budget of $n$ samples for the design, we can define D-optimal design as the best allocation $\{n_{x}\}_{x\in\mathcal{X}}\in\mathbb{N}$ of the $n$ samples such that $\log\det\left({\sum_{x\in\mathcal{X}}n_{x}xx^{\top}}\right)$ is maximized and $\sum_{x\in\mathcal{X}}n_{x}=n$. Finding a D-optimal design with a budget of $n$ is a challenging combinatorial optimization problem. However, there is a key property of the $\log\det$ function that enables an efficient approximation scheme. Namely, for a set $\mathcal{X}\subset\mathbb{R}^{d}$, the set function $S\subset\mathcal{X}\to\log\det\left({\lambda I+\sum_{x\in S}xx^{\top}}\right)$ is submodular if $\lambda$ is greater than one. Submodularity is a property describing decreasing additional benefit, known as diminishing returns. Fortunately, a submodular set function can be approximately maximized using a greedy algorithm, [Nemhauser et al., 1978]. Therefore, using Algorithm 3, we know that if $t_{\operatorname{stop}}$ and $t_{\operatorname{stop}}^{\prime}$ are two consecutive update times, then the design matrix satisfies: $\log\det\left({V_{t_{\operatorname{stop}}^{\prime}}}\right)\geq(1-1/e)\max_{S\subset\mathcal{X}}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{\top}}\right)$.
@@ E.2 On the intractability of optimistic approaches
Optimism in the face of uncertainty is a widely used principle for regret minimization and exploration in reinforcement learning. In the bandit setting, optimistic algorithms can be applied directly and yield minimax-optimal regret bounds [Auer, 2002]. However, in more general settings–such as linear bandits or MDPs–these approaches often lead to intractable optimization problems.
Optimism for regret minimization. For regret minimization in RLHF, an optimistic algorithm would choose the policy $\pi_{t}$ to maximize the upper confidence bound on the reward difference relative to a comparator policy $\pi_{t}^{\prime}$, typically the previous policy:
@@
where $\tilde{\theta}_{t}\in\mathcal{E}_{t}$. The cumulative regret can then be bounded using standard elliptical potential arguments (see Lemma A.5).
Optimism for pure exploration. Optimistic algorithms are also successful for minimizing the suboptimality gap of the final policy. In this case, algorithms aim to maximize the information gain by selecting either a policy pair $(\pi_{t},\pi_{t}^{\prime})$ or a single policy $\pi_{t}$ relative to a reference policy $\pi_{t}^{\prime}$ to maximize the exploration bonus:
$\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}.$
These formulations are appealing due to their theoretical guarantees. However, not much can be said about the computational efficiency of the subsequent optimization problems.
Challenge of optimizing the exploration bonus. As shown by Efroni et al. [2021], optimizing these exploration bonuses is intractable even in tabular MDPs with a known transition law. Specifically, maximizing the bonus reduces to the following problem over the set of discounted occupancy measures:
$\max_{\mu\in\mathcal{M}}\langle r_{\theta},\mu\rangle+\beta_{t}\left\lVert\mathbb{E}_{(s,a)\sim\mu}\phi(s,a)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}},$
where :=$\mathcal{M}:=\left\{{\mu^{\pi}:\mu^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s,a_{t}=a),\pi:\mathcal{S}\to\Delta_{\mathcal{A}}}\right\}$ is the set of valid occupancy measures induced by stationary policies [Puterman, 2014]. The norm term is strictly convex in $\mu$ due to the strict convexity of the norm and the linearity of expectations. As a result, this is a convex maximization problem, which is NP-hard in general [Atamtürk and Gómez, 2017].
These intractability results highlight a major limitation of optimistic algorithms. Although they yield strong statistical guarantees, their practical implementation in general MDPs remains computationally challenging. This motivates exploring alternative approaches, such as randomized exploration, which can provide almost optimal theoretical guarantees while being computationally tractable.
@@ Appendix F Experiments
All experiments run on Isaac Lab’s unmodified Isaac-Cartpole-v0 environment using the default PPO configuration. We train over 30 RLHF iterations, using 30 steps of PPO at each iteration, and training is repeated for 10 independent seeds. For the randomized exploration, we set $\beta_{t}=0.01+\max(1,\log t)$ and $\lambda=1$, and for lazy updates we set $C=1$. At each RLHF iteration we compare 5 independently sampled trajectories of horizon $H=150$. For the maximum likelihood estimation we perform 500 Adam steps (batch size $64$, $\ell_{2}$ penalty $\lambda=10^{-1}$). Experiments were executed on a single machine equipped with an Intel i9-14900KS CPU and an NVIDIA RTX 4090 GPU; completing 30 RLHF iterations required approximately 2 min 50 s.
Pure exploration
@@
Our results for the pure exploration setting with $\alpha=0$ are shown in Figure 2 and 3 below. In Figure 2, we see that all three versions of RPO achieve performance competitive to RL with the ground truth reward [^4] Note that the RL baseline makes $20024\times 4096\times 50=4^{\prime}915^{\prime}$ 200 queries to the ground truth reward, whereas RPO uses at most 150 binary preference queries., but LRPO-OD needs the least preference queries. Moreover, in Figure 3 we see that the performance during training is poor, i.e. the regret is large, which is to be expected due to the pure exploration scheme.
@@
Figure 2: Comparison of RLHF algorithms in terms of (a) the last iterate ground truth reward $V_{\theta^{*}}^{\hat{\pi}}$ (estimated from samples) and (b) number of preference queries performed. In particular, for the choice $\alpha=0$, we compare RPO (green, Algorithm 1) with it’s lazy versions LRPO and LRPO-OD (orange & blue). Here, LRPO and LRPO-OD refer to Algorithm 2 without and with optimal design subroutine. The error bars indicate the 0.2 and 0.8 quantiles, across 10 independent runs. The dashed blue line indicates the mean reward achieved by PPO with the ground truth parameter $\theta^{*}$.
@@
Figure 3: Comparing the ground truth rewards $V_{\theta^{*}}^{\pi_{t}}$ (estimated from samples) of RLHF algorithms for pure exploration (i.e. $\alpha=0$) during training, using the same color codes as in Figure 2.

### v2

@@ E.1 Background on optimal experimental design
The Kiefer-Wolfowitz theorem, see [kiefer1960equivalence], shows that a D-optimal design also ensures that $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\left({\sum_{x\in\mathcal{X}}\pi(x)xx^{\top}}\right)^{-1}}^{2}=d$. A direct consequence is that $\mathcal{X}$ is a subset of the ellipsoid $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$. The Kiefer-Wolfowitz theorem can be interpreted by saying that the set $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$ is the minimum volume ellipsoid containing $\mathcal{X}$, see [lattimore2020bandit, Theorem 21.1].
In the case where we have a budget of $n$ samples for the design, we can define D-optimal design as the best allocation $\{n_{x}\}_{x\in\mathcal{X}}\in\mathbb{N}$ of the $n$ samples such that $\log\det\left({\sum_{x\in\mathcal{X}}n_{x}xx^{\top}}\right)$ is maximized and $\sum_{x\in\mathcal{X}}n_{x}=n$. Finding a D-optimal design with a budget of $n$ is a challenging combinatorial optimization problem. However, there is a key property of the $\log\det$ function that enables an efficient approximation scheme. Namely, for a set $\mathcal{X}\subset\mathbb{R}^{d}$, the set function $S\subset\mathcal{X}\to\log\det\left({\lambda I+\sum_{x\in S}xx^{\top}}\right)$ is submodular if $\lambda$ is greater than one. Submodularity is a property describing decreasing additional benefit, known as diminishing returns. Fortunately, a submodular set function can be approximately maximized using a greedy algorithm, [nemhauser1978analysis]. Therefore, using Algorithm 4, we know that if $t_{\operatorname{stop}}$ and $t_{\operatorname{stop}}^{\prime}$ are two consecutive update times, then the design matrix satisfies: $\log\det\left({V_{t_{\operatorname{stop}}^{\prime}}}\right)\geq(1-1/e)\max_{S\subset\mathcal{X}}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{\top}}\right)$.
@@ E.2 On the intractability of optimistic approaches
Optimism for regret minimization Optimism in the face of uncertainty is a widely used principle for regret minimization in reinforcement learning. In the bandit setting, optimistic algorithms can be applied directly and yield minimax-optimal regret bounds [auer2002using]. For regret minimization in RLHF, an optimistic algorithm would choose the policy $\pi_{t}$ to maximize the upper confidence bound on the reward difference relative to a comparator policy $\pi_{t}^{\prime}$:
@@
where $\tilde{\theta}_{t}\in\mathcal{E}_{t}$. The cumulative regret can then be bounded using standard elliptical potential arguments (see Lemma D.4).
Optimism for preference-free exploration As is evident from the proof of Theorem 3.2.2, algorithms that solely maximize an exploration bonus of the form
$\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\quad\text{or}\quad\mathbb{E}_{\tau\sim\mathbb{P}_{\pi},\tau^{\prime}\sim\mathbb{P}_{\pi_{t}^{\prime}}}\left\lVert\phi(\tau)-\phi(\tau^{\prime})\right\rVert^{p}_{V_{t}^{-1}},p=1,2,$
are also effective for preference-free exploration.
Challenge of optimizing the exploration bonus The key problem of the above approaches is that for trajectory-level feedback – preferences or reward – they lead to optimization problems over policies that cannot be framed in terms of state-action rewards. efroni2021reinforcement therefore conjecture that exactly solving these problems is intractable even in tabular settings. This stands in contrast to standard reinforcement learning with state–action feedback, where optimistic algorithms, and approaches based on optimal design [wagenmaker2022beyond, mutny2023active], involve (or can be reduced to) maximization of bonuses of the form
$\mathbb{E}_{(s,a)\sim\mu_{\pi}}\left\lVert\phi(s,a)\right\rVert^{p}_{V^{-1}},p=1,2,$
which are linear in the occupancy measure $\mu_{\pi}$ and can be maximized via dynamic programming.
These computational challenges associated with exploration under trajectory-level feedback motivate alternative approaches, such as randomized exploration, which can provide near-optimal theoretical guarantees while remaining computationally tractable.
@@ Appendix F Experiments
F.1 Gridworld environment
As illustrated in Figure 3 below, the gridworld consists of 36 grid cells and the initial state lies in the center. The agent can choose the actions up, down, left, right, and will deterministically move in that direction or stay if it hits a boundary. The reward features are one-hot features for the six boundary states in blue, and the ground truth reward is $0.5$ for two of these states as indicated in Figure 3. Moreover, rewards are discounted with $\gamma=0.9$. Our gridworld implementation builds on the code by schlaginhaufen2023identifiability.
Figure 3: Illustration of the gridworld environment.
F.2 Cartpole environment
@@
All experiments run on Isaac Lab’s unmodified Isaac-Cartpole-v0 environment using the default PPO configuration. As reward features, we use the pre-defined reward terms. For cartpole these are: 1) Alive term: equal to 1 for all non-terminal states; 2) Termination term: equal to 1 for terminal states. A state is terminal if the cart goes out of bounds; 3) Goal tracking term: absolute value of pole angle measured from the upright position; 4) Cart velocity term: absolute value of cart velocity; 5) Joint velocity term: absolute value of pole angular velocity. However, our implementation supports any Isaac Lab manager‑based tasks.
@@
We train over 30 RLHF iterations, using 30 steps of PPO at each iteration, and training is repeated for 20 independent seeds. For the randomized exploration, we set $\beta_{t}=0.001+0.1\max(1,\log t)$ and $\lambda=1$, and for lazy updates we set $C=0.5$. At each RLHF iteration we compare 100 independently sampled trajectories. For the maximum likelihood estimation we perform 50 Adam steps (batch size $64$, $\ell_{2}$ penalty $\lambda=10^{-1}$). Experiments were executed on a single machine equipped with an Intel i9-14900KS CPU and an NVIDIA RTX 4090 GPU; completing 30 RLHF iterations required approximately 2 min 50 s.
Preference-free exploration
@@
Our results for the preference-free exploration algorithm RPO-Explore are shown in Figure 4 and 5 below. In Figure 4, we see that all three versions of RPO-Explore achieve performance competitive to RL with the ground truth reward [^4] Note that the RL baseline makes $24\times 4096\times 50=4^{\prime}915^{\prime}200$ queries to the ground truth reward, whereas RPO uses at most 150 binary preference queries., but RPO-OD-Explore needs the least preference queries. Moreover, in Figure 5 we see that the performance during training is poor, i.e. the regret is large, which is to be expected due to the pure exploration scheme.
@@
Figure 4: Comparison of RLHF algorithms in terms of (a) the last iterate ground truth reward $V_{\theta^{*}}^{\hat{\pi}}$ (estimated from samples) and (b) number of preference queries performed. In particular, we compare RPO-Explore (green, Algorithm 1) with its lazy versions LRPO-Explore and LRPO-OD-Explore (orange & blue). Here, LRPO-Explore and LRPO-OD-Explore refer to Algorithm 3 without and with optimal design subroutine. The error bars indicate the 0.2 and 0.8 quantiles, across 10 independent runs. The dashed blue line indicates the mean reward achieved by PPO with the ground truth parameter $\theta^{*}$.
@@
Figure 5: Comparing the ground truth rewards $V_{\theta^{*}}^{\pi_{t}}$ (estimated from samples) of RLHF algorithms for reward-free exploration during training, using the same color codes as in Figure 4.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide our theoretical claims in Sections 3 and 4, we also provide our empirical results in Section 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We address the limitations of our work in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide our assumptions in Section 2 and our proofs in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experimental environment will be publicly available, and all experimental details can be found in Appendix F.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code used for our experiment with the supplementary files.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details are provided in Appendix F.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Our experiment section provides shaded areas for the standard deviation of the performance curves over 10 independent training runs. Additional details can be found therein.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See appendix F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We don’t see any conflict with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not foresee any negative societal impact for our algorithms.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release models or datasets in this paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We are citing and crediting the Isaac Lab creators for their RL environment.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Instructions for reproducing the experiments are provided in the readme.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our research is not involving any experiments with humans.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our research is not involving any experiments with humans.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: We only use LLMs to improve the writing quality.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.